It is known that the relative permeability of iron is 100. If the mean atomic dipole moment is equal to 2microB . Find the magnets of magnetization M ?


1
Expert's answer
2021-05-18T11:11:05-0400

"M=\frac{B}{\mu \mu_0}=\frac{2\cdot 10^{-6}}{100\cdot 4\pi\cdot 10^{-7}}=16~\frac{mA}{m}."